OpenStudy (anonymous):

Write the expression 4^8 / 4^2 as a single power

OpenStudy (fibonaccichick666):

so if you were multiplying like bases, what do you do with the exponents?

OpenStudy (mathstudent55):

Multiplication of powers with the same base (add the exponents): \(\large a^m \times a^n = a^{m + n}\) Division of powers with the same base (subtract the exponents): \(\large \dfrac{a^m}{a^n} = a^{m - n}\)

OpenStudy (fibonaccichick666):

@dianegrace